Minutes of the Trigger Operation Meeting (May 31, 2005)
- Trigger has been running smoothly during this period of May 24-May 31, 2005.
- Last Wednesday (May 25), the damage to the fibers caused by a gopher was seen in TSF plots (see HN posting for more details). This is an example for running with both the production DCZ and parasitic DCT systems come in useful for identifying where the problem could originate from.
- Dan is working on getting the M* efficiency plot for mu-pairs in FastMon and OPR MON.
- Some more plots for EMT will be in FastMon. This will be ready to put in JAS and fastMon .ps archive on Thursday or Friday of this week.
- One bad TSF board was sent to Scott in UK last week. We have 2 more boards which should be fixed (soon?) in Manchester.
- Ming showed some validation plots for TSF data.
- Ming provided the final version of the LED info sheet which is also linked in L1 Trigger web page. A hard copy of this is posted next to the TSF crate at IR2.
- Dan will work on the validation check for DCH-DCT calibration so that we can compare drift chamber hits with what Trigger sees.
Data Quality and Monitoring
- Su Dong updated the web page for L1 FastMon tasks. A summer student from Stanford will help with some of these tasks.
- New FastMon plots (including EMT and all ODF crate damage monitoring) will be installed by this week.
This page is created by Selina Li (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.